OpenStudy (anonymous):

THE EQUATION OF ACURVE IS Y+2X+8/X^2 iii)Show that the normal to the curve at the point (-2,-2) intersects the x- axis at the point (-10,0)

OpenStudy (anonymous):

sry guys the equation is y=2x+8/x^2

OpenStudy (anonymous):

\[(2x+8)/x^2\] Like this?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

the normal might be the second derivative if i remember this correctly

OpenStudy (amistre64):

4 (x+12)/x^4 is what the wolf spits out; see if that works for the slope of the line between the 2 points to test

OpenStudy (anonymous):

Wasnt the normal the negative reciprocal of the first derivative?

OpenStudy (amistre64):

-1/4 is the slope between the points: 4(-2+12)/16 = 10/4 = 5/2 bummer yeah, the perp slope lol

OpenStudy (amistre64):

i musta been doing calc3 vectors

OpenStudy (anonymous):

XD

OpenStudy (amistre64):

-(2x+8)/x^3; x=-2 -(-4+8)/-8 = 1/4 so slope is -4 perhaps?

OpenStudy (amistre64):

-10,0 2 , 2 ------- -8,2; slope = -1/4 hmmm

OpenStudy (anonymous):

the slope is -1/4 u good :)

OpenStudy (amistre64):

i cant get the slopes to match up, we sure there aint a typo, in yours or mine?

OpenStudy (amistre64):

other than 1/4 spose to be 1/2 the perps to -2 :)

OpenStudy (amistre64):

2 + 8/x^2 downs to: 0 -16/x^3 at x=-2 we get -16/-8 = 2, perps to -1/2 is locser

OpenStudy (amistre64):

forgot an "x" lol 2x + 8x^-2; 2 - 16x-3; 2 + 2 = 4; which perps tp -1/4 got it lol

OpenStudy (anonymous):

i don`t get it how did we arrive to -1/4?

OpenStudy (amistre64):

which part?

OpenStudy (amistre64):

equation; first derivative = slope at any point; normal is perp slope; perp slope of m = -1/m

OpenStudy (anonymous):

amistre=amazing XD could you explain how you did it to me too :)

OpenStudy (amistre64):

the eq was:\[2x+8x^{-2}\]to begin with

OpenStudy (amistre64):

once you determine the perp sloped derivative :) compare it with the slope of the line between the given points

OpenStudy (anonymous):

thats smart :)

OpenStudy (amistre64):

thnx, and good luck ;)

OpenStudy (anonymous):

thank u next tym :)